The weight in pounds p of an olbject with a mass m of 25 kilograms is given by the equation p/m = 2.2. How many pounds does the object weigh?
Answer by stanbon(75887)

The weight in pounds p of an object with a mass m of 25 kilograms is given by the equation p/m = 2.2. How many pounds does the object weigh?
-------------------------------------------
p/m = 2.2
p/25 = 2.2
p = 25*2.2
p = 55 lbs.
==============================
Cheers,
Stan H.
========